### v1

@@ Abstract
Surgical phase recognition has gained significant attention due to its potential to offer solutions to numerous demands of the modern operating room. However, most existing methods concentrate on minimally invasive surgery (MIS), leaving surgical phase recognition for open surgery understudied. This discrepancy is primarily attributed to the scarcity of publicly available open surgery video datasets for surgical phase recognition. To address this issue, we introduce a new egocentric open surgery video dataset for phase recognition, named EgoSurgery-Phase. This dataset comprises 15 hours of real open surgery videos spanning 9 distinct surgical phases all captured using an egocentric camera attached to the surgeon’s head. In addition to video, the EgoSurgery-Phase offers eye gaze. As far as we know, it is the first real open surgery video dataset for surgical phase recognition publicly available. Furthermore, inspired by the notable success of masked autoencoders (MAEs) in video understanding tasks (e.g., action recognition), we propose a gaze-guided masked autoencoder (GGMAE). Considering the regions where surgeons’ gaze focuses are often critical for surgical phase recognition (e.g., surgical field), in our GGMAE, the gaze information acts as an empirical semantic richness prior to guiding the masking process, promoting better attention to semantically rich spatial regions. GGMAE significantly improves the previous state-of-the-art recognition method ($6.4\%$ in Jaccard) and the masked autoencoder-based method ($3.1\%$ in Jaccard) on EgoSurgery-Phase. The dataset will be released at https://github.com/Fujiry0/EgoSurgery.
@@ 1 Introduction
To tackle this issue, we introduce EgoSurgery-Phase, the first large-scale egocentric open surgery video dataset for phase recognition. 21 videos of procedures of 10 distinct surgical types with a total duration of 15 hours conducted by 8 surgeons are collected and annotated into 9 phases. The videos have been meticulously pre-processed for de-identification. EgoSurgery-Phase offers a rich collection of video content capturing diverse interactions among individuals (e.g., surgeons, assistant surgeons, anesthesiologists, perfusionists, and nurses), varied operative settings, and various lighting conditions. Moreover, in addition to video, EgoSurgery-Phase provides eye gaze data.
@@ 2.1 Dataset collection
Following the dataset collection protocol proposed in prior research [3], which focused on constructing datasets for surgical tool detection in open surgery videos, we gathered 21 open surgery videos utilizing Tobii cameras attached to the surgeon’s head. The recording of patient videos received ethical approval from the Keio University School of Medicine Ethics Committee, and written informed consent was obtained from all patients or their guardians. Our dataset encompasses 10 distinct types of surgeries, performed by 8 different surgeons.
The 21 videos were recorded at a frame rate of 25 fps and a resolution of $1920\times 1080$ pixels. Video durations vary between 28 and 234 minutes, reflecting the diversity in type and complexities of surgery. In total, 28 hours of surgical footage were captured. Unlike videos of minimally invasive surgery (MIS), open surgery videos are more likely to contain personally identifiable information (PII) such as the faces of patients, assistant surgeons, and nurses. To address privacy concerns, we subsampled the videos to 0.5 fps and anonymized the patient’s face through blurring. In addition, we exclude frames containing other PII. After these pre-processing steps, the average duration of the videos becomes 46 minutes, resulting in a total duration of 15 hours, thereby yielding a large-scale dataset of high quality. In addition to video, EgoSurgery-Phase provides eye gaze.
@@ 2.2 Dataset annotation, statistics and data split
Expert surgeons perform the annotations based on their clinical experience and domain knowledge. The 21 pre-processed videos of open surgery are manually annotated into 9 phases: Disinfection, Design, Anesthesia, Incision, Dissection, Hemostasis, Irrigation, Closure, and Dressing. Samples are shown in Fig. 1. In total, $27,694$ frames are manually annotated. The sample distribution is shown in Fig.3. It reveals a notable class imbalance. We use $14$ videos for the training set, $2$ videos for the validation set, and $5$ videos for the test set.
@@ 4.3 Phase recognition performance comparison
Comparison with masked autoencoder-based methods. After being pre-trained with the proposed GGMAE framework, the model exhibits significant performance improvements compared to the model trained from scratch ($6\%$ improvement in the Jaccard index). We then compare current state-of-the-art MAE-based methods, namely VideoMAE [15] and VideoMAEV2 [19]. Additionally, we evaluate our approach against SurgMAE [6], which first demonstrates the effectiveness of MAEs in the surgical domain. The performance of all methods is summarized in Table 2. Employing the same backbone and training schema, GGMAE surpasses VideoMAE by $4.1\%$ and VideoMAEV2 by $3.1\%$ and SurgMAE by $6.1\%$ in terms of Jaccard index.

### v3

@@ Abstract
Surgical phase recognition has gained significant attention due to its potential to offer solutions to numerous demands of the modern operating room. However, most existing methods concentrate on minimally invasive surgery (MIS), leaving surgical phase recognition for open surgery understudied. This discrepancy is primarily attributed to the scarcity of publicly available open surgery video datasets for surgical phase recognition. To address this issue, we introduce a new egocentric open surgery video dataset for phase recognition, named Egosurgery-Phase. This dataset comprises 15 hours of real open surgery videos spanning 9 distinct surgical phases all captured using an egocentric camera attached to the surgeon’s head. In addition to video, the Egosurgery-Phase offers eye gaze. As far as we know, it is the first real open surgery video dataset for surgical phase recognition publicly available. Furthermore, inspired by the notable success of masked autoencoders (MAEs) in video understanding tasks (e.g., action recognition), we propose a gaze-guided masked autoencoder (GGMAE). Considering the regions where surgeons’ gaze focuses are often critical for surgical phase recognition (e.g., surgical field), in our GGMAE, the gaze information acts as an empirical semantic richness prior to guiding the masking process, promoting better attention to semantically rich spatial regions. GGMAE significantly improves the previous state-of-the-art recognition method ($6.4\%$ in Jaccard) and the masked autoencoder-based method ($3.1\%$ in Jaccard) on Egosurgery-Phase. The dataset is released at project page.
@@ 1 Introduction
To tackle this issue, we introduce Egosurgery-Phase, the first large-scale egocentric open surgery video dataset for phase recognition. 20 videos of procedures of 10 distinct surgical types with a total duration of 15 hours conducted by 8 surgeons are collected and annotated into 9 phases. The videos have been meticulously pre-processed for de-identification. EgoSurgery-Phase offers a rich collection of video content capturing diverse interactions among individuals (e.g., surgeons, assistant surgeons, anesthesiologists, perfusionists, and nurses), varied operative settings, and various lighting conditions. Moreover, in addition to video, EgoSurgery-Phase provides eye gaze data.
@@ 2.1 Dataset collection
Following the dataset collection protocol proposed in prior research [3], which focused on constructing datasets for surgical tool detection in open surgery videos, we gathered 20 open surgery videos utilizing Tobii cameras attached to the surgeon’s head. The recording of patient videos received ethical approval from the Keio University School of Medicine Ethics Committee, and written informed consent was obtained from all patients or their guardians. Our dataset encompasses 10 distinct types of surgeries, performed by 8 different surgeons.
The 20 videos were recorded at a frame rate of 25 fps and a resolution of $1920\times 1080$ pixels. Video durations vary between 28 and 234 minutes, reflecting the diversity in type and complexities of surgery. In total, 28 hours of surgical footage were captured. Unlike videos of minimally invasive surgery (MIS), open surgery videos are more likely to contain personally identifiable information (PII) such as the faces of patients, assistant surgeons, and nurses. To address privacy concerns, we subsampled the videos to 0.5 fps and anonymized the patient’s face through blurring. In addition, we exclude frames containing other PII. After these pre-processing steps, the average duration of the videos becomes 46 minutes, resulting in a total duration of 15 hours, thereby yielding a large-scale dataset of high quality. In addition to video, EgoSurgery-Phase provides eye gaze.
@@ 2.2 Dataset annotation, statistics and data split
Expert surgeons perform the annotations based on their clinical experience and domain knowledge. The 20 pre-processed videos of open surgery are manually annotated into 9 phases: Disinfection, Design, Anesthesia, Incision, Dissection, Hemostasis, Irrigation, Closure, and Dressing. Samples are shown in Fig. 1. In total, $27,694$ frames are manually annotated. The sample distribution is shown in Fig.3. It reveals a notable class imbalance. We use $14$ videos for the training set, $2$ videos for the validation set, and $5$ videos for the test set.
@@ 4.3 Phase recognition performance comparison
Comparison with masked autoencoder-based methods. After being pre-trained with the proposed GGMAE framework, the model exhibits significant performance improvements compared to the model trained from scratch ($6\%$ improvement in the Jaccard index). We then compare current state-of-the-art MAE-based methods, namely VideoMAE and VideoMAEV2. Additionally, we evaluate our approach against SurgMAE, which first demonstrates the effectiveness of MAEs in the surgical domain. The performance of all methods is summarized in Table 2. Employing the same backbone and training schema, GGMAE surpasses VideoMAE by $4.1\%$ and VideoMAEV2 by $3.1\%$ and SurgMAE by $6.1\%$ in terms of Jaccard index.
@@ 5.0.1 Acknowledgements
5.0.2 Disclosure of Interests.
The authors have no competing interests to declare that are relevant to the content of this article.